Absorbance of the sample solution is 0.152AU. The three main uncertainty sources of absorbance in photometric analysis are repeatability, drift of the photometer and uncertainty due to possible interfering effects. The respective uncertainty components were estimated by the analyst as follows: Repeatability standard deviation is 0.005AU, uncertainty due to possible drift is ± 0.001AU and uncertainty due to the possible interfering effects is ± 2% of the absorbance value.

Please calculate the combined standard uncertainty of absorbance. Please provide 4 decimal places after the decimal point.
1
Expert's answer
2020-04-19T15:30:51-0400

Given : Absorbance of the sample solution is "0.152AU"

Repeatability standard deviation = "0.005 AU \implies e_1=0.005 AU"

Uncertainty due to possible drift "=± 0.001AU \implies e_2= 0.001AU"

Uncertainty due to the possible interfering effects is "± 2\%" of the absorbance value.

"\implies e_3=0.02*(0.152)=0.00304=0.0030AU" (rounded off upto four decimal places).


where "e_1,e_2 , e_3" are errors due to uncertainty in repeatability, drift of the photometer and due to possible interfering effects respectively.

Thus, the combined standard uncertainty of absorbance "=e_1+e_2+e_3=0.0050+0.0010+0.0030=0.0090 AU" ---(Answer)